# CoVR: Learning Composed Video Retrieval from Web Video Captions

### v3

@@ Appendix D Qualitative analysis
In this section, we provide examples of caption filtering (Section D.1), qualitative comparison between different MTG approaches (Section D.2), qualitative examples of our WebVid-CoVR triplets (Section D.3), samples from our manual test set annotation process (Section D.4), qualitative CoVR results on WebVid-CoVR-Test (Section D.5) and CoIR results on CIRR (Section D.6).
D.1 Examples of filtered captions
@@
D.2 Qualitative comparison of MTG approaches
@@
D.3 Training triplet examples
Figures A.9A.10, and A.11 all show examples of triplets generated using our automatic dataset creation. These examples demonstrate the effectiveness of our approach in generating coherent modification texts for paired videos. This capability serves as a form of data augmentation and increasing the diversity in the training set. In Figure A.12, we show that the dataset is not composed by pairs only, as there are many captions that have many relations between them. Furthermore, in Figure A.13 we show cases where a single caption is associated with multiple videos. This scenario allows us to generate multiple triplets by leveraging the diverse visual content captured in different videos. The triplets shown in the aforemention figures exhibit a wide range of variations, encompassing different themes such as emotions, food, actions, camera edits, gender changes, and time of the day.
D.4 Manual test set annotation
@@
D.5 Qualitative CoVR results on WebVid-CoVR-Test
@@
D.6 Qualitative CoIR results on the CIRR benchmark

### v4

@@ E Qualitative analysis
In this section, we provide examples of caption filtering (Section E.1), qualitative comparison between different MTG approaches (Section E.2), qualitative examples of our WebVid-CoVR triplets (Section E.3), samples from our manual test set annotation process (Section E.4), qualitative CoVR results on WebVid-CoVR-Test (Section E.5) and CoIR results on CIRR (Section E.6).
E.1 Examples of filtered captions
@@
E.2 Qualitative comparison of MTG approaches
@@
E.3 Training triplet examples
Figures A.11A.12, and A.13 all show examples of triplets generated using our automatic dataset creation. These examples demonstrate the effectiveness of our approach in generating coherent modification texts for paired videos. This capability serves as a form of data augmentation and increasing the diversity in the training set. In Figure A.14, we show that the dataset is not composed by pairs only, as there are many captions that have many relations between them. Furthermore, in Figure A.15 we show cases where a single caption is associated with multiple videos. This scenario allows us to generate multiple triplets by leveraging the diverse visual content captured in different videos. The triplets shown in the aforemention figures exhibit a wide range of variations, encompassing different themes such as emotions, food, actions, camera edits, gender changes, and time of the day.
E.4 Manual test set annotation
@@
E.5 Qualitative CoVR results on WebVid-CoVR-Test
@@
E.6 Qualitative CoIR results on the CIRR benchmark